# MidiCaps - A large-scale MIDI dataset with text captions

### v1

@@ 5 Conclusion
We present the first large-scale open MIDI captioned dataset, MidiCaps. This dataset also includes a comprehensive set of musical features such as chord patterns, genre, and mood. To facilitate the development of this dataset, we have developed a MIDI captioning framework. This approach includes music feature extraction and summarization from MIDI and the synthesized audio, as well as the use of the Claude-3 LLM to generate the final captions using in-context learning. To evaluate the final dataset, we have conducted two subjective listening studies, which confirm that the captions are natural and indeed contain a text description of the musical features contained in the accompanying MIDI file. The resulting new MidiCaps dataset contains 168,407 MIDI files with descriptive text captions is available online555huggingface.co/datasets/amaai-lab/MidiCaps under a Creative Commons licence.

### v2

@@ 5 Conclusion
We present the first large-scale open MIDI captioned dataset, MidiCaps. This dataset also includes a comprehensive set of musical features such as chord patterns, genre, and mood. To facilitate the development of this dataset, we have developed a MIDI captioning framework. This approach includes music feature extraction and summarization from MIDI and the synthesized audio, as well as the use of the Claude-3 LLM to generate the final captions using in-context learning. To evaluate the final dataset, we have conducted two subjective listening studies, which confirm that the captions are natural and indeed contain a text description of the musical features contained in the accompanying MIDI file. The resulting new MidiCaps dataset contains 168,407 MIDI files with descriptive text captions and is available online555huggingface.co/datasets/amaai-lab/MidiCaps under a Creative Commons licence.